Grade 12th passMechanics

Estimate the number of breaths taken during an average human lifetime.

hilal shah
6 Years agoGrade 12th pass

1 Answer

Arun
ApprovedApproved Tutor Answer6 Years ago
Let's assume that average lifespan for normal human being is 70 years.
In 70 years there are 365 days in each year.
70 years = 70 × 365 days
In one day there are 24 hours.
70 years = 70 × 365 × 24 hours
In one hour there are 60 minutes.
70 years = 70 × 365 × 24 × 60 minutes
we estimate an average respiration rate of about 10 breaths per minute. ... (Given)
Thus, for 70 years number of breath are,
n = 70 × 365 × 24 × 60 × 10
n = 3.679 × 108